katherine:

Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form. y = x + 8 y = x − 8 y = −5x − 10 y = −5x + 20

katherine:

@Vocaloid @Bearclaws72

SourMunchkin7806:

A

katherine:

you're sure?

SourMunchkin7806:

yea lol...would you like me to explain

katherine:

would to like for yourself to explain?

katherine:

if so, Please, be my guest :)

SourMunchkin7806:

so y=x+8...look at the number pairs...we will use (2,10)...the number on the left would be X and the right would be Y...so if we use the equation y=x+8, and plug that into the number pair we get 10=2+8...which if my math is correct...is the right answer

SourMunchkin7806:

got it?

katherine:

yes

SourMunchkin7806:

ok welp then thats your answer lol

katherine:

awesome. ty :)